### v1

# Planet formation theory: an overview

Abstract
The standard model for planet formation is a bottom-up process in which the origin of rocky and gaseous planets can be traced back to the collision of micron-sized dust grains within the gas-rich environment of protoplanetary disks. Key milestones along the way include disk formation, grain growth, planetesimal formation, core growth, gas accretion, and planetary system evolution. I provide an introductory overview of planet formation, emphasizing the main ideas and reviewing current theoretical understanding. Many of the phases of planet formation have a well-developed physical understanding, though the complexity of the problem means that few can be quantitatively modeled with complete confidence. Transformative advances in disk imaging and exoplanet detection provide the first direct information on the initial conditions for planet formation, and have motivated new formation models that are faster, more efficient, and lead to a more diverse set of architectures than their Solar System inspired forebears. Much remains to be learned, and I close with a personal, incomplete list, of open problems.
1 Introduction
Since the work of Safronov (1972), planet formation has been understood to be a bottom-up process whose initial phases play out within the dust and gas of protoplanetary disks. In contemporary versions of the theory, microscopic grains collide and stick to form pebbles, which participate in collective processes that form km (and larger) scale planetesimals via self-gravity. A combination of gravitationally mediated planetesimal collisions, and aerodynamically assisted pebble accretion, drives growth of the planetesimals up to the scale of Earth masses. If disk gas remains present, it can be captured by sufficiently massive cores to form the primary atmospheres or envelopes of the Solar System’s ice- and gas-giants, and of sub-Neptune and larger exoplanets. Less massive planets acquire secondary atmospheres from outgassing. Substantial orbital changes (migration) commonly occur after planets have formed, due to energy and angular momentum exchange between planets, between planets and gaseous or planetesimal disks, and between planets and their stellar hosts or binary companions.
This review provides an elementary introduction to this standard model, and serves as an overview of the following Chapters which discuss planet formation theory in detail. I introduce the physical processes relevant for planet formation, and summarize how they fit together in our consensus understanding of how planets form. Although a plausible outline of planet formation theory has been known for decades, existing models are surely not the last word: despite the genuinely transformative advances in exoplanet detection and disk imaging there remain few direct observations of planet formation processes. It is conceivable that are fundamental gaps in our current theory. Even within the established paradigm, some phases are known to be poorly understood, and I conclude by listing some open questions that I hope to see resolved soon.
2 Disk formation and evolution
Although limited grain growth may occur during star formation prior to disk formation (Kwon et al. 2009; Marchand et al. 2023), the vast majority of planet formation processes occur within protoplanetary disks. Disks are flattened near-equilibrium fluid structures, whose basic physical properties are determined observationally to varying (mostly low) levels of precision. Typical stellar accretion rates scale roughly as $\dot{M}\sim 10^{-8}(M_{*}/M_{\odot})^{2}M_{\odot}{\rm yr}^{-1}$, with a measurement precision of $\sim 0.35\ {\rm dex}$ and a spread at fixed stellar mass $M_{*}$ of more than an order of magnitude (Manara et al. 2023). The disk lifetime is around 3 Myr (Haisch et al. 2001), though the absolute ages of Young Stellar Objects are poorly determined (Soderblom et al. 2014). The relative ages of young clusters are more reliable. Disk masses can be estimated starting from various observables, including the dust continuum flux (Ansdell et al. 2016; Tychoniec et al. 2020), molecular line emission in commonly observed sub-mm tracers (CO and its isotopologues, sometimes corrected using additional species such as N2H+; Trapman et al. 2022), HD emission (McClure et al. 2016), and (at the high mass end) rotation curves (Martire et al. 2024; Andrews et al. 2024). All of these methods have drawbacks – either from limited applicability or from possibly large uncertainties – such that a consensus empirical model for disk evolution remains elusive. Planet formation theorists often assume as initial conditions a disk gas surface density profile that scales with radius as $\Sigma\propto r^{-1}$, an initial gas to dust ratio of $10^{2}$, and a mass within 100 au of $10^{-3}-0.1\ M_{*}$, but this is more a matter of art and convention than anything based on solid data. Despite its well-known shortcomings, the Minimum Mass Solar Nebula (Weidenschilling 1977b; Hayashi 1981) also lives on in some computers.
Some important properties of protoplanetary disks follow directly from the fact that they are long-lived, and thus in near force balance. In the “vertical” direction (perpendicular to the disk mid-plane) there is hydrostatic equilibrium between the gradient of gas pressure $P$ and the vertical component of stellar gravity,
$\frac{{\rm d}P}{{\rm d}z}=-\Omega_{\rm K}^{2}\rho z,$ (1)
where $\rho$ is the density, $\Omega_{\rm K}$ the Keplerian angular velocity, and we have assumed that $z\ll r$, the radial distance to the star. For a vertically isothermal disk with sound speed $c_{s}$, $P=\rho c_{s}^{2}$ and the gas density declines as a gaussian,
$\rho=\rho_{0}\exp[-z^{2}/2h^{2}],$ (2)
where $\rho_{0}$ is the mid-plane density and $h$, the vertical scale height, is,
$h\equiv\frac{c_{s}}{\Omega_{\rm K}}.$ (3)
Except in cases of unusually high accretion rates, or when close to the star, the heating of protoplanetary disks occurs primarily from the interception and re-radiation of stellar luminosity (Kenyon and Hartmann 1987). Irradiation leads to a radial temperature profile approximately given by a power-law, $T\propto r^{-3/7}$ (Chiang and Goldreich 1997), and a disk that flares, such that $h/r$ increases with radius. Typically, $0.10.03<h/r<$ 0.1. In the radial direction, the force balance involves gravity, centrifugal force, and a pressure gradient,
$\Omega^{2}(r)r=\frac{GM_{*}}{r^{2}}+\frac{1}{\rho}\frac{{\rm d}P}{{\rm d}r}.$ (4)
The pressure gradient term is order $(h/r)^{2}$ smaller than the gravitational one, so to leading order $\Omega(r)=\Omega_{\rm K}$. The small correction – which usually but not always leads to sub-Keplerian rotation – is nonetheless extremely important for planet formation. The dynamics of dust particles (Whipple 1972; Weidenschilling 1977a), the formation of planetesimals (Youdin and Goodman 2005), and the interpretation of observed disk sub-structures (Andrews 2020) are all thought to hinge on this $\sim$1deviation from strict Keplerian motion.
Figure 1: An illustration of the main processes that lead to evolution of the gaseous component of protoplanetary disks.
In contrast to the structural questions sketched above, little can be deduced about the evolution of protoplanetary disks from elementary physical principles. (Near)-Keplerian disks have a specific angular momentum profile, $l(r)\propto r^{1/2}$, that increases with radius, so for gas to accrete angular momentum has to be lost. Broadly speaking three classes of physical processes can lead to disk evolution,
  • Infall of gas, which will add mass and mix angular momentum to a pre-existing disk. Infall is self-evidently important at sufficiently early times, when disks are forming and growing (Lin and Pringle 1990), but it could also be significant at late times if the density of residual gas in the disk environment is high enough (Throop and Bally 2008; Winter et al. 2024; Padoan et al. 2024).
  • Processes that redistribute angular momentum within the disk, while conserving disk mass (up to the fraction of gas that is physically accreted by the star). Lesur et al. (2023) review the long laundry list of possible processes, most of which are associated with linear instabilities of disk flow. Self-gravity (Toomre 1964; Kratter and Lodato 2016), the magnetorotational instability (Balbus and Hawley 1991) as (strongly) modified by non-ideal magnetohydrodynamic effects, and the Vertical Shear Instability (Nelson et al. 2013) are the leading suspects.
  • Loss of mass and / or angular momentum through disk winds. Photoevaporation (Hollenbach et al. 1994) – in which the surface of the disk is heated by high-energy photons to the point where it escapes in a thermal wind – is guaranteed to occur at some level (Alexander et al. 2014; Ercolano and Pascucci 2017). It is a pure mass-loss process. Magnetized winds (Blandford and Payne 1982), on the other hand, remove gas that has more angular momentum than the disk at the launch point, simultaneously dispersing the disk while driving accretion (Bai and Stone 2013a, b)
Figure 1 illustrates these basic possibilities. In addition to evolution, many of these processes can also plausibly generate rings, spirals, and other sub-structures in disks.
The state-of-the-art in direct numerical simulation (Xu and Kunz 2021; Mauxion et al. 2024) allows disks to be followed for about $10^{5}\ {\rm yr}$, roughly the duration of the Class 0 phase of Young Stellar Object evolution. Longer duration evolution, and studies that seek to compare theory against observed populations of disks, necessarily require simpler models. The minimal goal of one-dimensional models is to predict $\Sigma(r,t)$, though more complete treatments may also evolve the disk central temperature (this is essential for modeling disk outbursts; Bell and Lin 1994; Armitage et al. 2001; Zhu et al. 2009). The vertical magnetic flux threading the disk, $B_{z}(r,t)$, is another important dynamical quantity that in principle should be evolved in models that invoke magnetized winds or the magnetorotational instability. The simplest case is a disk that evolves solely due to local turbulent redistribution of angular momentum. The appropriate one-dimensional description is a diffusion equation (Lynden-Bell and Pringle 1974; Pringle 1981),
$\frac{\partial\Sigma}{\partial t}=\frac{3}{r}\frac{\partial}{\partial r}\left[r^{1/2}\frac{\partial}{\partial r}\left(\nu\Sigma r^{1/2}\right)\right],$ (5)
in which $\nu$ is an effective viscosity that is generated by turbulence. If $\nu$ is approximated by a time-independent power-law function of radius, this equation admits a similarity solution (a time-varying exponentially truncated power-law for the surface density) that has been widely used for disk modeling (e.g. Hartmann et al. 1998). More generally, the viscosity can be written in terms of the local scale height $h$ and disk sound speed $c_{s}$ as (Shakura and Sunyaev 1973),
$\nu=\alpha c_{s}h,$ (6)
with $\alpha$ being a dimensionless parameter that characterizes the efficiency of angular momentum transport. Wind-driven evolution can be modeled within an analogous framework, by adding sink and advective terms to the right hand side of equation (5) that represent the effects of mass and angular momentum loss on the underlying disk. The detailed form of the additional terms depends on the choice of wind model; Tabone et al. (2022) derive a generalization of the Lynden-Bell and Pringle (1974) solution for one choice that may be generic enough to be useful for population modeling. Some effects due to the infall of gas onto the disk are also amenable to a (numerical) one-dimensional description (Lin and Pringle 1990).
Determining the relative importance of these evolution processes, from either theoretical or observational arguments, is a major challenge. It has often been assumed that the ordering from top to bottom in Figure 1 maps onto a temporal sequence, in which disks form from infall on a time scale of $\sim 10^{5}\ {\rm yr}$, evolve initially as massive self-gravitating structures and then more slowly under the action of viscosity and winds, and are finally dispersed by photoevaporation. This remains a plausible paradigm. Observations of molecular line broadening (Flaherty et al. 2017), and of the thickness of dust layers (Villenave et al. 2022), imply low levels of turbulence in many (though not all; Flaherty et al. 2024) of the best-studied disks. The inferred values of $\alpha$ in the “weakly turbulent” disks – typically $\alpha\sim 10^{-4}$ or lower – are broadly consistent with strongly damped MHD turbulence (described as “dead zone” models; Gammie 1996) and weak purely hydrodynamic driving mechanisms. As a consequence, magnetized disk winds rather than turbulence are assumed to be dominant in many recent models. There remains room, however, for larger deviations from conventional wisdom, such as models in which infall is dynamically important at late as well as early times.
A primary result of ALMA (Atacama Large Millimeter/submillimeter Array) has been the discovery that protoplanetary disks typically show pronounced substructure, in the form of departures from the simplest model of an axisymmetric disk with a monotonically decreasing profile of intensity (Andrews 2020). Substructure can be detected most economically in dust continuum imaging, but is also seen in molecular line data that directly traces the gas. Rings, crescent-shaped features, and spirals are all observed, roughly in decreasing order of frequency. Substructure is also detected in scattered light images of disks (Avenhaus et al. 2018), most prominently around more massive stars. Scattered light substructure arises from small grains in the disk atmosphere, and is often harder to interpret than sub-mm images that trace regions closer to the mid-plane. Necessarily, all of the observed substructure involves orbital radii and spatial scales $>{\rm au}$. There are robust theoretical reasons for the existence of cavities and rings at sub-au scales, including from the interaction of a stellar magnetosphere with the disk (Zhu et al. 2024) and from the onset of stronger turbulence due to the magnetorotational instability at temperatures where the alkali metals are thermally ionized (Gammie 1996). These unobserved structures could be central to our understanding of close-in planet formation (Chatterjee and Tan 2014).
The proximate cause of currently observed substructure in dust continuum images is understood to be the aerodynamic concentration of particles in local gas pressure maxima (or, in weaker cases, local pressure enhancements over a smooth background). This process – often described imprecisely as “trapping” – was foreseen by Whipple (1972) (if you haven’t already, look at his amazingly prescient Figure 1). The pressure maxima can be in form of axisymmetric rings (Haghighipour and Boss 2003), as in the original Whipple proposal, vortices (Barge and Sommeria 1995), or spiral waves (Rice et al. 2004), matching the main classes of observed structures. Ring-like dust structures are also predicted to form without the need for underlying gas perturbations in the vicinity of condensation / sublimation fronts (Zhang et al. 2015), but this effect does not appear to be responsible for most of the substructure that is currently being seen.
While the immediate cause of dust substructure is reasonably clear, the same cannot be said for why gas disks develop pressure maxima in the first place. A popular hypothesis is that the source of substructure is gravitational perturbations from already-formed planets. Planets launch spiral density waves, and open gaps whose edges – which are approximately axisymmetric pressure maxima – can become unstable to vortex generation through the Rossby Wave Instability (Lovelace et al. 1999). Planet-disk interaction models generate disk substructure that is often a good match to what is actually observed (Zhang et al. 2018), and are surely responsible for some substructure. That said, simulations of disk turbulence and magnetized winds also commonly lead to the formation of rings (“zonal flows”; Johansen et al. 2009) and vortices, while self-gravity of course forms spirals. As a result, the interpretation of substructure is uncertain. For the purposes of this review, the key takeaway is that the observations of substructure have profound implications for planet formation irrespective of their interpretation. If ringed disks almost always host planets, that means that a substantial population of ice or gas giant planets is able to form surprisingly quickly at orbital separations that are as large as $\sim 50~{}{\rm au}$. If substructure is instead a byproduct of disk transport processes, then the observed concentration of solids that results modifies the evolution of all subsequent phases of planetary growth.
3 Aerodynamic and collisional evolution of solids
Dust inherited by the disk is assumed to follow the size distribution, $n(a)\propto a^{-3.5}$, that is inferred from extinction measurements in the interstellar medium (Mathis et al. 1977). Growth up to the scale of small macroscopic solids (“pebbles”, which might be anywhere between mm and meters in physical size) is controlled by aerodynamic and collisional processes. The physics of these processes is quite subtle, but most of the important results follow from two considerations:
  1. 1.
    The measured material properties of silicate and water ice aggregates imply that fragmentation occurs when collision velocities exceed a threshold value (nominally 1 ms-1, though possibly up to 10 ms-1 for ice in some circumstances). Collision speeds in a turbulent gas increase with particle size, so this material limit translates to a predicted maximum physical particle size, normally in the mm-cm range.
  2. 2.
    Solid particles are coupled to the gas via aerodynamic drag forces, but imperfectly so. Deviations from perfect coupling have multiple consequences, including vertical settling, radial drift, particle concentration in vortices, and small-scale clustering and collisions. Back-reaction of the aerodynamic forces on the gas, when the dust to gas density ratio is non-negligible, leads to the streaming instability. These effects depend not on the physical particle size, but on a dimensionless measure of “size” for aerodynamic purposes called the Stokes number.
Figure 2 illustrates some of these processes, which operate all the way from observable global scales down to microscopic scales on which dust particles are colliding.
Figure 2: Aerodynamic coupling between solid particles and disk gas controls the early phases of planet formation on multiple scales. On the smallest scale, it leads to clustering of solids (simulation from Chan et al. 2024) and collisions between individual particles (Ormel and Cuzzi 2007). On intermediate scales, it leads to the streaming instability (Youdin and Goodman 2005) and planetesimal formation (simulation from Simon et al. 2016). The Kuiper Belt contact binary Arrokoth (bottom right panel) is consistent with this formation channel. On large, observable scales, aerodynamics leads to vertical settling, radial drift, and particle concentration at pressure maxima.
Small dust grains in gaseous disks experience low velocity collisions due to Brownian motion (Dullemond and Dominik 2005), that lead to sticking and the formation of porous aggregates. Larger particles collide at higher velocities, due to coupling to small-scale turbulence or size-dependent drift, leading to compactification and, eventually, bouncing or fragmentation. A microscopic theory (“JKRS theory”, applied to the astrophysical context by Dominik and Tielens 1997) relates the collision outcome to the surface energies of the materials involved, and can be used to model the growth of aggregates via molecular dynamics techniques (e.g. Wada et al. 2009). The fiducial results for grain growth, however, come from laboratory experiments, usually performed in microgravity conditions (Blum and Wurm 2008). For silicate particles, considered to be representative of the inner disk interior to the snowline, net growth occurs for collisions with relative velocity $\Delta v<1\ {\rm ms}^{-1}$ (Blum and Wurm 2008). For water ice, collision experiments by Gundlach and Blum (2015) at $T\approx 100-200\ {\rm K}$ found growth for $\Delta v<10\ {\rm ms}^{-1}$. Measurements of water ice surface energies, however, are similar to values measured for silicates at temperatures $T<175\ {\rm K}$ (Musiolik and Wurm 2019). In JKRS theory this would imply that both silicate and ice aggregates cease to stick above similar (low) threshold collision speeds, except near the snowline where water ice unambiguously shows a higher threshold. There is evidently some uncertainty in this key quantity, and it is common in models to consider the sticking and fragmentation thresholds for ice as parameters and explore the effect of different values in the 1-10 ms-1 range.
Translating the experimental results for the sticking and fragmentation thresholds into a predicted maximum particle size $a_{\rm max}$ requires knowledge of what sets $\Delta v$ as a function of size $a$. In the most common situation, collisions at $a\sim a_{\rm max}$ are caused by aerodynamic coupling to small-scale turbulence (this process also leads to small-scale clustering of solids; Cuzzi et al. 2001). For particles of size $a_{1}$ and $a_{2}$, coupled aerodynamically to isotropic Kolmogorov turbulence, Ormel and Cuzzi (2007) provide the standard (in the astrophysical literature) calculation that gives $\Delta v(a_{1},a_{2})$. Adopting an $\alpha$ model for the turbulence, the peak collision speed is found to be,
$(\Delta v)_{\rm max}\sim\alpha^{1/2}c_{s},$ (7)
where $c_{s}$ is the sound speed. Noting that $c_{s}=h\Omega_{\rm K}$, and taking $\alpha=10^{-4}$ and $h/r=0.1$, this result implies that $(\Delta v)_{\rm max}>1\ {\rm ms}^{-1}$ wherever the orbital velocity exceeds 1 kms-1 (for a Solar mass star, out to almost 1000 au). One concludes that the material properties of particles often impose a strong constraint on growth. Detailed calculations (including radial drift, which is usually a stronger limiter at large radii) show that the maximum size of particles that form from pairwise collisional evolution is perhaps 0.1 mm at 100 au, increasing to around 1 cm at 3 au (Birnstiel 2024).
Aerodynamic processes in disks depend on the stopping time of particles. A particle of mass $m$, moving at speed $\Delta v$ relative to the gas, feels an aerodynamic drag force $F_{\rm drag}$ and slows down on a time scale,
$t_{\rm stop}=\frac{m\Delta v}{|F_{\rm drag}|}.$ (8)
It turns out to be very useful to multiply $t_{\rm stop}$ by a characteristic frequency and make the stopping time dimensionless. In isotropic turbulence (relevant to the grain-grain collisions discussed above) the appropriate characteristic frequency is that of eddies, but for most other applications the orbital angular frequency is the right choice. This is called the Stokes number (or, the dimensionless stopping or friction time),
$\tau_{\rm S}\equiv t_{\rm stop}\Omega.$ (9)
The form of the drag force depends on particle size (Weidenschilling 1977a). For $a<\lambda$, the mean free path of molecules in the gas, we are in the Epstein regime, and for a spherical particle,
${\bf F}_{\rm drag}=-\frac{4\pi}{3}\rho a^{2}v_{\rm th}\Delta{\bf v}.$ (10)
Here, $\rho$ is the gas density and $v_{\rm th}$ its thermal speed (roughly equal to the sound speed). With this drag law, the stopping time is independent of $\Delta v$ and the dimensionless Stokes number is fixed by the size and composition of the particle. In the inner disk, the gas density is high enough that particles near the fragmentation limit have $a>\lambda$. This is the Stokes regime, in which the gas flows as a fluid around the particle and $F_{\rm drag}\propto(\Delta v)^{2}$. Cheng (2009) provides a parameterization of the drag force that is valid across both regimes.
In the vertical direction, particles settle toward the disk midplane at approximately their terminal velocity. Settling is opposed by turbulent diffusion. In the tracer limit (i.e. assuming that the volume density of dust is negligible compared to the gas), and assuming that the magnitude of turbulent diffusion (in the vertical direction) equals the viscosity, $D=\nu$, the equilibrium particle distribution is a gaussian (Dubrulle et al. 1995). The scale height of the particles, $h_{\rm d}$, is related to that of the gas through (Birnstiel 2024),
$\frac{h_{\rm d}}{h}=\sqrt{\frac{\alpha}{\tau_{\rm S}+\alpha}}.$ (11)
This allows observations of the thickness of dust disks to be used to constrain disk turbulence. The assumption that $D=\nu$ is quite crude, and there can be substantial deviations from this simple formula in disks where the turbulence is driven by the Vertical Shear Instability (Stoll and Kley 2016; Pfeil et al. 2024; Fukuhara and Okuzumi 2024) or by self-gravity (Baehr and Zhu 2021).
In the radial direction, aerodynamic effects lead to drift of particles relative to the gas. For $\tau_{\rm S}\ll 1$, the strong aerodynamic coupling means that dust orbits at almost the same speed as the gas, but, unlike the gas, it does not feel the pressure gradient. The unbalanced radial force leads to drift in the direction of higher gas pressure, which in a smooth disk would be inward. The general expression for the drift speed, valid not just in the trace particle limit, was derived by Nakagawa et al. (1986). If the gas orbits at speed, $v_{\phi}=(1-\eta)v_{\rm K}$, then dust with a dust to gas ratio $\epsilon=\rho_{\rm dust}/\rho$ drifts radially at speed (e.g. Birnstiel 2024),
$v_{\rm r,dust}=-\frac{2\eta v_{\rm K}}{\tau_{\rm S}+(1+\epsilon)^{2}\tau_{\rm S}^{-1}}.$ (12)
Even though $\eta\sim(h/r)^{2}$ is small, typically only $\sim{\rm few}\times 10^{-3}$, the rate of the resulting drift is highly consequential. At 1 au, the time scale for drift of particles with $\tau_{\rm S}=1$ would be as short as of the order of a thousand years. Particles with this Stokes number would be of the order of meters in size, leading to the moniker “meter-sized barrier” for this aerodynamic constraint on planetesimal formation. (Material physics limits on sticking already make the growth of meter-sized boulders improbable, but the meter-sized barrier provides an additional, independent, argument against planetesimal formation by pairwise coagulation.) More importantly in a modern context, particles with physical sizes of mm-cm and Stokes numbers $\tau_{\rm S}\sim 10^{-2}$, that can readily form given our best understanding of coagulation physics, are predicted to suffer substantial drift even in the outer disk on scales of 10-100 au.
In disks with dynamically insignificant substructure, radial drift is inward. It leads to an effective limit on particle size: at any radius, particles cannot grow to a size for which their growth time scale exceeds the radial drift time scale (Birnstiel et al. 2012). Defining $\gamma=|\partial\ln P/\partial\ln r|$, the maximum particle size in drift-limited growth is (Birnstiel 2024),
$a_{\rm max,drift}=\frac{2\Sigma_{\rm dust}}{\pi\rho_{\rm s}\gamma}\left(\frac{h}{r}\right)^{-2},$ (13)
where $\Sigma_{\rm dust}(r)$ is the dust surface density and $\rho_{\rm s}$ is the material density. This limit can be compared to the limit imposed by fragmentation (Birnstiel 2024),
$a_{\rm max,frag}=\frac{2\Sigma}{3\pi\rho_{\rm s}\alpha}\left(\frac{\Delta v_{\rm frag}}{c_{s}}\right)^{2}.$ (14)
The drift limit is usually the most stringent in the outer disk, while fragmentation limits growth closer to the star.
Turbulent diffusion in the radial direction can be important both for the evolution of trace gas species, and for the evolution of dust populations that are simultaneously subject to radial drift (Morfill and Voelk 1984; Clarke and Pringle 1988). Tominaga et al. (2019) derive formulae for dust diffusion in disks that conserve angular momentum and which can be used to model dust concentration processes and possible instabilities. An important consequence of diffusion is the qualitative change it makes to how dust evolves in the vicinity of gas pressure maxima. Ignoring diffusion, an axisymmetric pressure maximum – whether it forms spontaneously or at a planetary gas edge – is efficient at trapping aerodynamically drifting solids, filtering all but the smallest particles that are swept inward by the mean flow (Rice et al. 2006). With diffusion, it is always possible to find a steady state solution in which the turbulent flux overcomes the aerodynamic trapping effect, such that pressure maxima temporarily detain but do not permanently imprison solids (Ward 2009; Zhu et al. 2012b).
Birnstiel (2024) reviews dust growth in protoplanetary disk in much more detail than can be covered here. DustPy (Stammler and Birnstiel 2022) is open source software that can be used to calculate how solids grow and drift within disks.
4 Planetesimal formation
Planetesimals are primordial bodies whose sizes, although not precisely defined, are usually considered to be in the range between 100 m and a few hundred km. The growth of planetesimal-scale bodies is a key milestone in planet formation because it marks the double onset of gravity’s dominance. Above some critical size, probably around a few hundred meters (Benz and Asphaug 1999), self-gravity rather than material strength controls the outcome of physical collisions. Above some other, rather larger physical size, gravitational forces become more important than aerodynamic ones in the dynamics of a population of bodies orbiting within the disk.
Theories on how planetesimals form have shifted markedly over time. Goldreich and Ward (1973), and independently Safronov (1972), suggested that dust settling proceeds until the mid-plane dust density exceeds the Roche density, which at orbital radius $r$ is,
$\rho_{\rm Roche}\sim\frac{M_{*}}{r^{3}}.$ (15)
If $\rho_{\rm dust}>\rho_{\rm Roche}$, the solid material can collapse gravitationally to form planetesimals. Although there are circumstances where this process may operate, it is no longer considered to be a generally viable planetesimal formation mechanism. Typically, vertical shear will generate enough turbulence to halt settling before the critical density is reached, even in an intrinsically non-turbulent disk (Cuzzi et al. 1993). The other obvious mechanism for planetesimal formation is a continuation of pairwise sticking collisions from pebble to planetesimal scales (Weidenschilling and Cuzzi 1993). Again, while this model cannot be entirely ruled out, it is largely unviable given current understanding of fragmentation and radial drift processes.
The current working hypothesis for planetesimal formation is that the streaming instability (Youdin and Goodman 2005) generates particle clumps that are dense enough to collapse gravitationally to planetesimals. The streaming instability is a local, linear instability of the Nakagawa et al. (1986) solution for equilibrium radial drift. It requires consideration of the momentum feedback particles exert on the gas, and a non-zero drift velocity (hence, it cannot formally be operative at the peak of a pressure maximum). Youdin and Goodman (2005) show that the streaming instability has non-zero (though sometimes small) growth rates for almost all combinations of Stokes number and dust to gas ratio, and it should therefore be generically present in protoplanetary disks. Jacquet et al. (2011) and Magnan et al. (2024) discuss the rather subtle physical mechanism of the streaming instability, which is the most important (for planet formation) of a broader class of resonant drag instabilities (Squire and Hopkins 2018). It should be noted that the Nakagawa et al. (1986) equilibrium studied by Youdin and Goodman (2005) differs in multiple ways from both real disks and from the setups for most streaming instability simulations. The Nakagawa et al. (1986) model is vertically unstratified (vertical gravity is generally not negligible), does not include turbulent diffusion (which is present if there is any instrinsic disk turbulence and which is needed to establish a stratified particle layer equilibrium), and is based on a fluid description of the particles (this is only a good approximation for $\tau_{\rm S}\ll 1$, and is not used in particle-based simulations). Subsequent linear calculations have included the effects of a particle size distribution (Krapp et al. 2019; Paardekooper et al. 2020; Zhu and Yang 2021) and turbulent diffusion (Chen and Lin 2020; Umurhan et al. 2020; Gerbig et al. 2024), which act to reduce the growth rates and restrict the range of parameters over which linear instability is present.
Numerical simulations, starting with the pioneering work of Youdin and Johansen (2007) and Johansen et al. (2007), have demonstrated a streaming-initiated pathway to planetesimal formation. It is illustrated in Figure 2 (using simulations from Simon et al. 2016). From uniform initial conditions, the streaming instability generates largely axisymmetric filaments of particles on small scales of the order of $\sim 0.1h$. Such structures may have observable effects on disk emission (Scardoni et al. 2021), but they are not presently able to be directly imaged. (They are not the rings seen in ALMA disk images.) Within the filaments dense clumps form, and given favorable conditions these clumps can exceed the Roche density and collapse gravitationally. The expected outcome of collapse, which cannot be followed to material density with current simulation techniques, is some combination of single planetesimals, binaries (Nesvorný et al. 2010, 2021), or small clusters of planetesimals.
Simon et al. (2022) review planetesimal formation physics, and the large simulation literature. One important set of results concerns the threshold dust to gas ratio needed for planetesimal formation, defined via a metallicity $Z=\Sigma_{\rm dust}/\Sigma$. Carrera et al. (2015) found that the critical metallicity for “prompt” (i.e. basically dynamical time scale) planetesimal formation, as a function of the Stokes number, follows a U-shaped curve with a minimum near $\tau_{\rm S}\simeq 0.1$. Subsequent revisions to the critical curve (Yang et al. 2017; Li and Youdin 2021; Lim et al. 2024b) show that for a standard value of the pressure gradient, $Z_{\rm crit}$ is below 0.01 – the fiducial value of the global dust to gas ratio in disks – for $0.01\leq\tau_{\rm S}\leq 1$. These thresholds apply to the most favorable scenario of negligible intrinsic disk turbulence and a monodisperse particle size distribution. When turbulence is present, the criterion for planetesimal formation is approximately that the settled particle layer has a mid-plane density $\rho_{\rm dust}/\rho>1$ (Gole et al. 2020; Lim et al. 2024a).
A second important simulation result is the predicted mass function of planetesimals formed from the streaming instability. A large number of studies agree that streaming-initiated collapse leads to a top-heavy initial planetesimal mass function, in which most of the total mass is in the largest objects (Johansen et al. 2007; Simon et al. 2016; Schäfer et al. 2017; Li et al. 2019; Schäfer et al. 2024). The absolute size of the largest planetesimals is of course a function of the disk conditions, but for reasonable parameters it is often hundreds of kilometers. This first principles prediction is large compared to what was assumed in most formation-agnostic planetary growth models in the past, and has substantial downstream implications both for the Solar System (“asteroids were born big”; Morbidelli et al. 2009) and for planet formation generally. In particular, 100+ km planetesimals experience much less aerodynamic drag than 100 m-1 km planetesimals, and when their orbits are excited by gravitational perturbations the population will saturate at a higher velocity dispersion. This reduces the rate at which protoplanets grow by planetesimal accretion, so that, counter-intuituvely, being able to rapidly make large primordial planetesimals does not necessarily imply faster overall planet formation time scales.
There is no analytic expectation for the form of the initial planetesimal mass function, but several simple functions (an exponentially tapered power-law, a broken power-law, etc) provide fits to numerical results that can be compared to Solar System populations of small bodies. The Kuiper Belt is widely considered to be the most primordial surviving population and offers the greatest potential for a test, though even there careful consideration has to be given to the extent of collisional evolution (Morbidelli et al. 2021; Benavidez et al. 2022). Other possible tests of the predictions of planetesimal formation models include the inclination distribution of Kuiper Belt binaries (Nesvorný et al. 2019), and the detailed morphology of the single small Kuiper Belt Object imaged by the New Horizons mission (Arrokoth; McKinnon et al. 2020). The results of these tests are consistent with streaming instability simulations.
Given the limitations of both the Solar System data and of current numerical simulations, tests of the streaming hypothesis for planetesimal formation are not decisive. Other proposals for planetesimal formation include collisional growth of porous icy aggregates (Okuzumi et al. 2012), and various non-streaming processes that can concentrate solids to above the threshold for gravitational collapse while avoiding the barriers that hinder the Goldreich-Ward mechanism. Examples include secular gravitational instability (Ward 1976; Youdin 2011; Takahashi and Inutsuka 2014), and particle concentration in spiral arms (Rice et al. 2004; Baehr et al. 2022; Longarini et al. 2023) or vortices (Lyra et al. 2024). Secular gravitational instability is interesting, in part, because the favorable range in Stokes number may be smaller than for streaming instability, while models that invoke gravitational collapse of particle clumps in disk substructures tend to produce extremely massive “planetesimals” that would more commonly be considered to be planetary embryos or dwarf planets.
5 Formation of rocky planets, cores, and giants
Once planetesimals have formed, their initial growth to become protoplanets, and ultimately either rocky planets or the cores of ice or gas giants, occurs via two physically distinct mechanisms:
  1. 1.
    Planetesimal accretion. The plantesimals physically collide, due to gravitational interactions, at low enough velocity to grow. The rate of collisions is determined, through the physics of gravitational focusing, by the velocity dispersion of the planetesimals. To a first approximation, one imagines the protoplanet growing from a local planetesimal reservoir that does not drift radially.
  2. 2.
    Pebble accretion. While the gas disk is still present, a fraction of appropriately aerodynamically coupled pebbles (typically, $10^{-3}\leq\tau_{\rm S}\leq 1$) that enter the protoplanet’s Hill sphere are gravitationally captured and accreted via a settling flow. The protoplanet accretes a fraction of the solids that drift past its orbit, with an efficiency that depends on its mass, on the Stokes number, and on the disk properties.
Figure 3 illustrates the three ways these mechanisms could combine for an overall scenario of planetary growth. The classical (pre-2010) model considers only planetesimal growth. This must be wrong at some level – as we observe roughly pebble-sized solids in disks – but planetesimal-driven growth could still be dominant at some times or locations in the disk. The other extreme is pure pebble growth. This is conceivable, but requires that at least some very massive planetesimals (probably, more massive than one would expect from the streaming instability) form. The most popular model is a hybrid, in which the earliest phases of growth occur via planetesimal collisions, with pebble accretion contributing and in some regimes dominating once larger protoplanets have formed. In any of these scenarios: planetesimal, hybrid, or pebble, further evolution occurs due to orbital instabilities (planet-planet collisions) and due to the accretion of envelopes by sufficiently massive cores.
Figure 3: Rocky planets (including the Solar System’s terrestrial planets and low mass exoplanets), and the cores of more massive planets, grow from planetesimals by a combination of planetesimal and pebble accretion. The classical model for terrestrial planet formation invokes only planetesimal growth, while more recent “hybrid” models involve an initial phase of planetesimal growth that is followed by predominantly pebble accretion. If the most massive planetesimals are large enough, a pure pebble model for subsequent growth may also be viable.
The key concepts of planetesimal-driven growth can be illustrated with a toy model for planet growth within an initially dynamically cold planetesimal disk. Consider a protoplanet of mass $M$, radius $R$, and surface escape speed $v_{\rm esc}$, orbiting within a disk of planetesimals with surface density $\Sigma_{p}$ and velocity dispersion $\sigma$. The typical eccentricity and inclination of the planetesimals are $e\sim i\sim\sigma/v_{K}$, so the planetesimal disk has thickness $\sim\sigma/\Omega$. Simple collision rate arguments (e.g Lissauer 1993; Armitage 2020) give a growth rate,
$\frac{{\rm d}M}{{\rm d}t}=\frac{\sqrt{3}}{2}\Sigma_{p}\Omega\pi R^{2}\left(1+\frac{v_{\rm esc}^{2}}{\sigma^{2}}\right),$ (16)
that describes growth in the so-called “dispersion dominated” limit (there is also a shear-dominated limit when tidal effects are dominant). The enhancement of the physical collision cross-section due to the effect of gravitational focusing is captured by the term in parenthesis. It can vary by several orders of magnitude, such that growth in the classical picture is closely tied to the evolution of $\sigma$. There are two regimes,
  • Runaway growth (Greenberg et al. 1978). A small protoplanet does not affect the local planetesimal velocity dispersion (typically, $\sigma$ is determined by balancing excitation from planetesimal-planetesimal scattering encounters against aerodynamic damping). Assume that $\sigma={\rm const}\ll v_{\rm esc}$. Then for two bodies in the same disk environment, with masses $M_{1}>M_{2}$, equation (16) predicts that ${\rm d}(M_{1}/M_{2})/{\rm d}t\propto(M_{1}/M_{2})(M_{1}^{1/3}-M_{2}^{1/3})>0$. Growth amplifies any initially small mass differences.
  • Oligarchic growth (Kokubo and Ida 1998). More massive protoplanets excite the velocity dispersion of nearby planetesimals, slowing their own growth while less massive neighbors catch up. This leads to a population of planetary embryos growing at comparable rates.
Ignoring migration the outcome is a system of protoplanets on low-eccentricity orbits, which begin to collide among themselves when their mass becomes comparable to the mass in the sea of remaining planetesimals (Kenyon and Bromley 2006). This end-point of planetesimal growth has traditionally been used as the effective initial condition for N-body simulations of the final assembly phase of rocky planets. Although these planetesimal-only models are rather simple – arguably too simple – they are quite successful in reproducing the basic properties of the Solar System’s terrestrial planets (Raymond et al. 2009).
Pebble accretion (Ormel and Klahr 2010; Lambrechts and Johansen 2012) is a qualitatively different physical process for planetary growth. While gas is present, small solid particles that encounter the planet within its Hill sphere are accreted if the time scale on which they would settle toward the planet is shorter than the time scale of the encounter. The accretion rate is particularly simple to estimate in the limit where the thickness of the pebble layer is larger than the Hill sphere of the planet,
$r_{\rm Hill}=\left(\frac{M}{3M_{*}}\right)^{1/3}r.$ (17)
Within the Hill sphere, by definition, the planet’s gravity dominates over the tidal gravitational field of the star, and it makes sense to define a time scale on which a pebble would settle toward the planet. Equating the drag force (equation 8) to the gravitational force for a pebble encountering the planet at impact parameter $b$ and relative velocity $\delta v$, the terminal velocity with which settling occurs is,
$v_{\rm settle}=\frac{GMt_{\rm stop}}{b^{2}}.$ (18)
Applying the condition that the settling time scale $t_{\rm settle}=b/v_{\rm settle}$ equals the encounter time scale $t_{\rm enc}=b/\delta v$ we find the relation that $\delta v=GMt_{\rm stop}/b^{2}$, and the pebble accretion rate is,
$\dot{M}\sim\rho_{\rm pebble}b^{2}\delta v\sim GM\rho_{\rm pebble}t_{\rm stop},$ (19)
where $\rho_{\rm pebble}$ is the density of the pebble “fluid”. This is an approximation for the pebble accretion rate in what is described as the three dimensional regime. Similar estimates can be derived in the two dimensional limit, where the pebble layer is thinner than the Hill sphere. In this regime, the accretion rate varies depending on whether the encounter velocity is due to disk shear, the headwind experienced by the pebbles as they drift, or a non-zero planetary eccentricity. Detailed expressions are given in Liu and Ormel (2018) and Ormel and Liu (2018). Generalizations to the case of a polydisperse size distribution of solids are given in Lyra et al. (2023).
Pebble accretion has several distinctive differences from planetesimal accretion. There is no regime of pebble accretion that exhibits runaway growth. It is also fairly strictly bounded in planet mass. Pebble accretion is inefficient below a pebble initiation mass (Visser and Ormel 2016), and for typical parameters would only be able to directly grow planetesimals if their initial sizes are rather large (at least several hundred km). Pebble accretion is also likely to be inefficient above a pebble isolation mass (Lambrechts et al. 2014; Bitsch et al. 2018), typically of the order of $50M_{\oplus}$. Above this mass, a growing planet starts to open a gap in the disk, forming a local pressure maximum that can halt the radial drift of solids.
Planetesimal accretion is severely limited at large orbital radii by the tendency of more massive bodies to scatter, rather than accrete, planetesimals (this is the origin of the long-known difficulty in forming Uranus and Neptune in situ from planetesimals; Levison and Stewart 2001). Pebble accretion, conversely, tends to work better at moderately large orbital radii – solids at the drift or fragmentation limit have more favorable aerodynamic properties, and there is no reduction in accretion efficiency as the ratio of Hill sphere size to planetary radius increases. Outside perhaps 5-10 au, pebble accretion is presumptively the dominant mechanism for planetary growth (e.g. Drazkowska et al. 2023). More general statements as to the relative role of planetesimals and pebbles in planet formation are, however, hazardous. Planetesimal and pebble accretion are distinct processes that cannot be easily compared, and while pebble accretion rates can be very high this is, in part, a consequence of the rapid radial drift of solids which cannot be sustained indefinitely. Which process dominates in the formation of observed planetary systems can only be considered in the context of a global theoretical model for planet formation (e.g. Bitsch et al. 2015), or by appealing to observational cosmochemical constraints that are only available for the Solar System. Johansen et al. (2021) argue for a large pebble contribution in terrestrial planet formation, but the division between the two processes in a hybrid Solar System model remains the subject of vigorous debate.
Accretion of the primary atmospheres or envelopes of massive planets occurs (in the consensus view) via the process of core accretion (Mizuno 1980), an umbrella term that encompasses the classical planetesimal-only model for Jupiter and Saturn’s formation (Pollack et al. 1996), along with more recent models that can accommodate the existence of the dilute core inferred for Jupiter from JUNO data (Wahl et al. 2017; Helled et al. 2022), or a primary role for pebble accretion (Lambrechts et al. 2014). The key insight of core accretion is the predicted existence of a sharp increase in the gas accretion rate when the envelope mass $M_{\rm env}$ exceeds a critical mass, that is approximately set at $M_{\rm env}\sim M_{\rm core}$ by the onset of self-gravity in the envelope (Rafikov 2006). Below the critical mass, low mass gaseous protoplanets grow relatively slowly, at a rate that is determined by the growth of the core through pebble and planetesimal accretion, and by the time scale on which the envelope can lose thermal energy and contract (Ikoma et al. 2000). If the gas disk dissipates during this initial phase of core accretion, the expected outcome is a planet that may resemble the Solar System’s ice giants, or extrasolar sub-Neptunes. If, instead, the critical mass can be exceeded during the disk lifetime, the planet enters a phase of accelerated gas accretion (this is known as “runaway growth”, although it is still a quasi-hydrostatic evolution) and the outcome is a gas giant with $M_{\rm env}\gg M_{\rm core}$. The minimum core mass that can lead to gas giant formation is primarily a function of the opacity in the envelope, and is roughly $5-10\ M_{\oplus}$ (Piso et al. 2015) for grain size distributions that arise in coagulation-fragmentation equilibrium calculations. Smaller critical core values are possible if envelopes are effectively grain-free (Hori and Ikoma 2010). In the other direction, substantially more massive cores can avoid entering runaway given sufficient heating from planetesimal accretion.
Recycling of gas between the protoplanetary disk and the outer envelope of a growing planet affects the thermodynamics of the envelope, and may therefore alter the time scale for achieving runaway (Ormel et al. 2015). The importance of this process depends upon the depth the recycled gas is able to reach within the envelope (Savignac and Lee 2024). Accurate modeling of recycling is a substantial challenge: the flows themselves are intrinsically three-dimensional and can only be captured fully in radiation hydrodynamic simulations, while their impact on core accretion is felt only over long time scales that can only be followed in one dimensional models. Current work suggests that the effect of recycling on the formation of Jupiter itself may be minimal (Zhu et al. 2021), but that for close-in (0.1 au) cores there could be as much as an order of magnitude delay in the onset of runaway growth (Bailey and Zhu 2024).
Figure 4: Illustration of a possible structure for a circumplanetary disk surrounding a massive planet. Gas is supplied to circumplanetary disks through out-of-plane flows that circularize at a small fraction of the planet’s Hill sphere. The inner disk accretes, while the outer disk may be approximately static (if it is tidally truncated) or decreting (if tidal truncation is inefficient). The disk close to the planetary surface terminates in an equatorial boundary layer if the planet is weakly magnetized, while strongly magnetized planets accrete via a magnetosphere that is analogous to accretion in T Tauri stars. Circumplanetary disks may be detected via dust continuum or molecular line emission, or via accretion signatures.
Following runaway, growth of giant planets to Jovian or super-Jovian masses is ultimately limited by the ability of the planet to accrete from across the gap that opens within the protoplanetary disk. Circumplanetary disks, fed by out-of-plane flows from the protoplanetary disk (Tanigawa et al. 2012), may form during this final phase of accretion. Circumplanetary disk formation requires not only a sufficiently large planet mass relative to the thermal mass (roughly, a mass such that $r_{\rm Hill}>h$; e.g. Sagynbayeva et al. 2024), but also favorable thermodynamic conditions that allow fast enough cooling on sub-Hill scales (Ayliffe and Bate 2009; Szulágyi et al. 2016; Fung et al. 2019). Multi-fluid radiation hydrodynamic simulations suggest that cooling on a time scale that is an order of magnitude shorter than the orbital time scale is necessary for gas to form a near-Keplerian disk on scales of $r_{\rm Hill}/3$ (Krapp et al. 2024). Observationally, the detection and characterization of circumplanetary disks is in its infancy. PDS 70c, a young massive exoplanet in a system whose protoplanetary disk has a transition disk morphology, shows sub-mm emission at the location of the planet that is consistent with circumplanetary disk expectations (Isella et al. 2019; Benisty et al. 2021). As illustrated in Figure 4, continuum emission is just one of many potential tracers of circumplanetary disks, which may also be observable in molecular (kinematic) or atomic line emission, or even through the production of jets (Gressel et al. 2013; Zhu 2015).
The consensus view is that the Solar System’s planets, and at least the vast majority of extrasolar planets, formed via core accretion (which, to be clear, is a broad framework for planet formation rather than a precisely specified model). It is possible that some giant planets, especially those approaching the $13\ M_{J}$ threshold that conventionally demarcates planets from brown dwarfs, instead formed via disk instability (Boss 1997). The formation of protostellar disks may well lead to initially self-gravitating disks (Lin and Pringle 1990; Xu and Kunz 2021), which have been observed to fragment into multiple stellar systems (Tobin et al. 2016). The physics of self-gravitating disks is reviewed by Kratter and Lodato (2016). Such disks fragment if the cooling time scale is sufficiently short (of the order of the orbital time scale; Gammie 2001; Rice et al. 2005), conditions that occur only at large ($\sim 50$ au) radii (Clarke 2009; Rafikov 2009). What are currently the most realistic radiation hydrodynamics simulations, by Xu et al. (2024), find that the distribution of initial fragment masses in this regime peaks at about $20\ M_{J}$. Further accretion in what remains a gas-rich environment post-fragmentation would likely increase fragment masses further (Kratter et al. 2010b). (This fate could only be averted if fragments migrate inward rapidly and are partially tidally stripped; Nayakshin 2010; Baruteau et al. 2011; Zhu et al. 2012a). Protostellar disks can also be driven to fragment by sufficiently rapid infall (Kratter et al. 2010a), but this also leads to characteristic masses that are typically in the brown dwarf or low-mass star regime.
6 Stability and planetary system evolution
Prescient work by Goldreich and Tremaine (1980) and Fernandez and Ip (1984) identified physical processes that can lead to the large-scale orbital migration of planets that interact gravitationally with gas or planetesimal disks. Their ideas, along with tidal and purely dynamical processes that can also reshape planetary orbits, received renewed interest with the discovery of the first exoplanets (Mayor and Queloz 1995), whose close-in and (separately) often eccentric orbits are inconsistent with our accepted understanding of in situ formation.
Planet-disk interactions, reviewed recently by Paardekooper et al. (2020), are one broad class of processes that lead to planetary system evolution. The most intensively studied case involves resonant interactions that allow the exchange of angular momentum with the protoplanetary disk, and migration of the planetary orbit. There are two sub-cases:
  • Gas at interior and exterior Lindblad resonances is excited by the planetary potential (Goldreich and Tremaine 1979), forming density waves that manifest as one-armed spirals in the disk (Ogilvie and Lubow 2002). It is by now well-understood. Key features of the Lindblad torque are that its strength is largely independent of disk structure (in the sense that how the waves dissipate is immaterial; Meyer-Vernet and Sicardy 1987), and that the radial pressure gradient generates a generic asymmetry between the summed inner and outer torques (Ward 1997). These properties mean that – although other torques have more complex behavior – the total torque on a planet from its interaction with a gaseous disk can only vanish coincidentally, or at special locations, known as migration traps (Hasegawa and Pudritz 2011). At planet masses low enough that the perturbation to the disk surface density profile is small (the “Type I” regime), the Lindblad-only contribution to migration is almost always inward.
  • Gas in the co-orbital region exchanges angular momentum with the planet as it executes horseshoe turns in the frame co-rotating with the planet. Asymmetries in the behavior of the turns in the regions that lead / trail the planet, due to radial gradients in entropy and vortensity, lead to a net torque. Complexity arise because, in the absence of viscosity or wind-driven radial motion, the streamlines in the horseshoe region are closed, and would rapidly adjust to null out (“saturate”) the torque. Simulations show that adding the co-orbital torque to the Lindblad torque can lead to outward migration, but that this result depends not just on the disk radial structure, but also on how energy and angular momentum are transported (Paardekooper et al. 2010, 2011).
More recently, significant non-resonant sources of migration torque have been identified. A luminous planet generates perturbations in temperature and density in its vicinity, that are asymmetric on either side of the orbit because (yet again) of the radial pressure gradient. This leads to a thermal torque, which has an established linear theory (Masset 2017). Besides altering the migration rate (in the sense of modestly broadening the range of parameters for which outward migration occurs; Guilera et al. 2019; Baumann and Bitsch 2020), an interesting property of thermal torques is their propensity to excite eccentricity and inclination (Fromenteau and Masset 2019; Cornejo et al. 2023). Another relatively new area of investigation is the role of pebble torques, that arise from the scattering of aerodynamically coupled particles that drift past a planet (Benítez-Llambay and Pessah 2018; Chrenko et al. 2024). These can also give rise to enhanced outward migration.
The coupled response of the planet-disk system to migration torques has traditionally been studied in the limit where angular momentum transport in the disk is due to turbulence, and the strength of that turbulence is moderately high ($\alpha\sim 10^{-3}-10^{-2}$). In this case, migration occurs in two regimes. Low mass planets do not open a gap, and migrate in the “Type I” regime due to the intrinsic asymmetry between inner and outer Lindblad resonance locations (Ward 1997). High mass planets open a gap, and their migration in the “Type II” regime is then coupled to the evolution of the disk (Lin and Papaloizou 1986). Whether a planet opens a gap, and hence what regime of migration applies, depends on a simple function of the planet-to-star mass ratio, disk scale height, and $\alpha$-parameter (Kanagawa et al. 2015). The accuracy of simplified (one dimensional) time-dependent models for Type II migration remains debated (Duffell et al. 2014; Scardoni et al. 2020; Kanagawa and Tanaka 2020), but it is established that the fastest migration typically occurs in the Type I regime just prior to gap opening (Ward 1997; Bate et al. 2003). In many planet formation applications, the inferred migration time scale is comparable to or shorter than the gas disk lifetime.
In low-viscosity disks, the picture is more complex. Planet-disk interaction in this regime commonly forms vortices through the Rossby Wave Instability (Lovelace et al. 1999), which can themselves migrate and interact with the planet. McNally et al. (2019) show how the migration regime varies in a two-dimensional plane, with the planet mass and disk viscosity as the controlling parameters. Particularly interesting is the case of a disk with a low turbulent viscosity, but efficient wind angular momentum loss. Wind angular momentum loss is essentially a surface rather than a volume process, so a wind-driven disk is expected to drive a rapid radial flow through planetary gaps that will affect migration torques (McNally et al. 2020; Wu and Chen 2025; Hu et al. 2024). Using three dimensional non-ideal magnetohydrodynamic simulations, Wafflard-Fernandez and Lesur (2023) find that magnetic flux concentration in the gap region leads to an asymmetric gap profile that leads to outward migration of massive planets.
In the linear, Type I regime, the disk torque on low mass planets is expected to be largely independent of the multiplicity of the planetary system. The migration of the system will, however, be affected if there is angular momentum exchange due to mean-motion resonances between the planets. Resonant capture (Goldreich and Peale 1966; Borderies and Goldreich 1984) occurs when there is convergent migration into a mean motion resonance, with a probability that depends on the migration speed, eccentricity, and strength of perturbations (due, in the case of migrating planets, to gravitational coupling to disk turbulence) (Quillen 2006; Adams et al. 2008; Batygin 2015). For giant planets in the Type II regime, overlap of the gaps created by the planets means that the Lindblad torques acting on them are no longer independent. Under the right conditions, two giant planets, that individually would each migrate inward, migrate outward as a coupled resonant pair (Masset and Snellgrove 2001). These rather subtle multi-planet effects are of central importance to the “breaking the chains” model for close-in Kepler systems (Izidoro et al. 2017), for modern version of the Nice model for the outer Solar System (Tsiganis et al. 2005; Morbidelli et al. 2007), and in the Grand Tack model for the inner Solar System (Walsh et al. 2011). Planetesimal scattering, subsequent to gas disk dispersal, provides a mechanism for breaking resonances, especially in the context of the Nice model. Given a sufficiently massive planetesimal disk, scattering may be important across a broader range of radii during planet formation (Minton and Levison 2014; Jinno et al. 2024).
Migration is a phenomenon that, once protoplanets exceed masses of a fraction of an Earth mass, impacts all phases of planet formation and planetary system evolution. It has a similar relation to planet formation as opacity does to stellar structure, in that it is critically important, complex, but in principle calculable from first principles to useful precision. There are only a few circumstances where completely ignoring migration is a defensible position.
After the dissipation of the protoplanetary disk, planets with atmospheres or envelopes can lose mass through photoevaporation (Owen and Wu 2013) and / or core-powered mass loss (Ginzburg et al. 2016). The immediate architecture emerging from the disk phase – during which planet-disk interactions typically (though not always) damp eccentricity and inclination – may also prove to be unstable over long time scales. The simplest model problem here is the evolution of a system of $N$ planets, initially on circular, coplanar orbits, under Newtonian point-mass gravity. For $N=2$, stability to close encounters between the planets (“Hill stability”) can be proven beyond a threshold separation (Gladman 1993). For $N\geq 3$, planetary systems are found through N-body integrations to be unstable, with a time scale to instability that is a steep function of the initial separation (measured in units of the planet mass to the 1/3 or 1/4 power; Chambers et al. 1996; Smith and Lissauer 2009; Obertas et al. 2017). These numerical findings are interpreted within a dynamical framework of chaos driven by the overlap of resonances (Lecar et al. 2001; Quillen 2011; Lammers et al. 2024). Resonant chain systems – arguably the most natural initial condition emerging from the disk phase – can of course have much longer lifetimes than similar planetary systems with non-resonant period ratios. Pichierri and Morbidelli (2020) and Goldberg et al. (2022) analyze the stability of this important special case.
The outcome of dynamical instability in a planetary system can be ejection of one or more of the planets, physical collisions, or excitation to a high eccentricity state that may collide or interact tidally with the host star. The parameter that controls the branching ratios between these outcomes is the ratio of the orbital velocity to the escape speed from the surface of the planet. High values of $v_{\rm K}/v_{\rm esc}$ (low mass planets, and / or small orbital radii) favor collisions, while low values lead to a predominance of ejections (Ford et al. 2001; Petrovich et al. 2014).
Planets that are sufficiently widely separated to be quasi-stable to scattering can still exhibit non-trivial secular evolution. Secular dynamics approximates conserves the angular momentum deficit (for a modern reference, see e.g. Laskar 1997),
${\rm AMD}\equiv\sum_{k}\Lambda_{k}\left(1-\cos i_{k}\sqrt{1-e_{k}^{2}}\right),$ (20)
where the sum is over planets with angular momenta $\Lambda_{k}$, semi-major axes $a_{k}$, eccentricities $e_{k}$, and inclinations (relative to the invariable plane) $i_{k}$. AMD conservation ensures that nothing interesting develops from initially near-circular, near-coplanar systems of widely separated planets (“AMD stability”; Laskar and Petit 2017). Given non-zero AMD, however (potentially arising from prior planet-planet scattering events), chaos driven by secular resonances leads to each planet eventually experiencing substantial excursions in AMD about an equipartition value (Wu and Lithwick 2011). A large AMD for an inner, relatively low mass planet, can lead to a high eccentricity that places the pericenter distance close enough to the star for tidal dissipation to become important. Qualitatively similar behavior is possible under Kozai-Lidov dynamics (Kozai 1962; Lidov 1962; Naoz et al. 2011), in which a planet exhibits cyclic variations in eccentricity due to secular perturbations from a distant, inclined companion. These secular processes can naturally produce high (and even retrograde) inclinations as well as high eccentricities (Wu and Murray 2003; Fabrycky and Tremaine 2007; Naoz et al. 2011), making them prime suspects in the formation of the misaligned exoplanets that have been detected via the Rossiter-Mclaughlin effect (Winn et al. 2010).
Dynamical processes are generally considered to play a dominant role in the excitation of giant planets to high eccentricities (Rasio and Ford 1996), and a substantial / dominant role (when combined with the action of tides) in the formation of hot Jupiters (Dawson and Johnson 2018). With the detection of planets around younger stars, and better age dating of older planetary systems, there is an exciting possibility to measure and interpret dynamically differences in planetary system architecture as a function of stellar age (Hamer and Schlaufman 2024; Dai et al. 2024). Nonetheless, thirty years on from the discovery of the first exoplanets, the relative importance of the many processes that could lead to hot Jupiter formation remains an open problem.
7 Some open questions
The other contributions to this section of the Handbook of Exoplanets review planet formation theory in much more detail than is possible here. In many areas there has been substantial recent progress. Theoretical advances, many inspired by problems thrown up by new Solar System, disk, and exoplanet observations, have led to a theory of planet formation that inherits many features of classical models while being more efficient and accommodating the formation of diverse final architectures. It is self-evident, however, that much remains to be understood. Open questions and problems that I would like to see resolved include, in no particular order:
  1. 1.
    What sets the balance between pebbles and planetesimals? Aerodynamically coupled particles, especially those with $\tau_{\rm S}\sim 0.1$, are valuable currency in contemporary theory, both for planetesimal formation via the streaming instability and for pebble accretion. Pebbles can only be spent once, however, and it is not clear how the balance is determined.
  2. 2.
    Uncertainty quantification of migration rates. Computing the rate of gas disk migration to a reliably known level of precision, across the range of planet masses and disk structures relevant to planet formation, would remove a major source of uncertainty. Code comparison efforts are a valuable step toward this goal.
  3. 3.
    Is planet formation smoothly distributed or radially concentrated at special disk locations? Protoplanetary disks contain ice lines, and probably other deterministic radial locations where the disk physics creates particle traps or planet migration null points. The basic paradigm for planet formation could yet be either smooth or ring-like.
  4. 4.
    Do disks evolve mostly due to turbulent or wind torques? Magnetized winds are currently favored, but even if the basic question is considered resolved, the strength and nature of the turbulence in primarily wind-driven disks is still of fundamental importance for planet formation.
  5. 5.
    Does planet formation overlap spatially or temporally with disk accretion outbursts, or with regions of active gravitational instability? Either of these processes – conventionally but not necessarily occurring mostly at early times – lead to a fundamentally different environment for growth.
  6. 6.
    What is the structure and evolution of gas and solids in circumplanetary disks? This is a physics-rich problem with a wide range of relevant spatial scales. Informative constraints from circumplanetary disk, and possibly exomoon observations, will be needed to inform theoretical work.
  7. 7.
    Does the composition of the upper levels of exoplanet atmospheres retain formation information? The division of abundant elements such as carbon and oxygen between condensed and gaseous phases is expected to vary with temperature, and thus radius, in the disk. This opens up the possibility of using atmospheric composition measurements from transmission spectroscopy as a new probe of planet formation (Öberg et al. 2011). Multiple confounding factors, in both planet formation and atmospheric evolution, might however frustrate such hopes.
  8. 8.
    How can theoretical models of planetary system evolution be tested against population-level data? Even when the theoretical model is well-specified and accurately calculable (e.g. pure N-body dynamics), and the data well-understood, inferring principled statistical constraints is often difficult. New statistical and Machine Learning approaches show promise.
  9. 9.
    What are the properties of planetesimals that form from the streaming instability, and from alternative planetesimal formation channels? Important quantities such as the initial mass function, and the distribution of binary inclinations, have already been derived from simulations of the streaming instability. How robust these predictions are, and to what extent they discriminate between models, is not completely clear: few simulations include collisional physics or have the resolution to follow collapse to anything close to material density.
  10. 10.
    How is the opacity, especially that from grains, determined within planetary envelopes? The efficiency of cooling, determined largely by the envelope opacity, controls the outcome of core accretion, and hence the predicted populations of sub-Neptunes, ice, and gas giants.
  11. 11.
    Are the conclusions of N-body, and joint statistical / N-body simulations of planet formation, robust? GPU-enabled N-body codes can evolve fully self-gravitating systems with substantially larger $N$ than was previously possible, allowing both a test of the robustness of existing results and the possibility of direct simulations of growth across a broader mass range.
  12. 12.
    Is the material physics of collisions well enough understood for planet formation applications? The experimental determinations of the threshold velocity for fragmentation of silicates and water ice, and more broadly the outcomes of different collisions between aggregates of various sizes, are critical inputs for planet formation theory. Any revision to the currently accepted paradigm, perhaps from experiments with more realistic materials, could have a major impact.
Acknowledgements.
My understanding of planet formation has been greatly informed by the work and ideas of students, postdocs, and other collaborators over many years. Thanks to all of you!